* Question Completion Status: QUESTION 6 Ang company that produces pleasure boats has decided to expand one of its lines. Theh
Answer #1

Solution:

Let,

FC = Fixed Cost

VC = Variable Cost Per Unit

FC VC Location L1 3.1 L2 7958 5879 11897 4.5 2.2 L3

Then,

Cross Over Point Between Two Locations = Absolute Difference Between Fixed Costs of Two Locations / Absolute Difference Between The Variable Cost Per Unit of Two Locations

Thus,

Coross Over Point Between L1 and L2 = | 7958 Coross Over Point Between L2 and L3 | 5879 Coross Over Point Between L1 and L3 1

(Rounded to the nearest whole number)

Hence,

Answer a:

The range of output for L1 resulting in the lowest total cost = 1485 to 4377

Answer b:

The range of output for L2 resulting in the lowest total cost = 0 to 1485

Answer c:

The range of output for L3 resulting in the lowest total cost = 4377 and More

Answer d:

TC = Total Cost = FC + (Q x VC)

Where Q = 4005 Boats

Total Cost For Location L1 = 7958 + ( 4005 X 3.1 (4005 x 3.1 ) = 20373.50 Total Cost For Location L2 = 5879 + (4005 x 4.5 ) =

Hence, the alternative location L1 yields the lowest total cost for the expected annual volume of 4005 boats

(Kindly raise an upvote for this answer, if you found it useful)